A balloon is floating around outside your window. The temperature outside is 29 ∘C , and the air pressure is 0.800 atm . Your neighbor, who released the balloon, tells you that he filled it with 4.50 moles of gas. What is the volume of gas inside this balloon?
Answer #1

Hi Dear Friend, here we can use simple ideal gas equation PV = nRT

R = gas constant = 0.082057 L.atm/mol.K; T = 29 0C = 29 + 273.15 = 302.15 K

V = nRT / P = [4.50 mol x 0.082057 L.atm/mol.K x 302.15 K] / 0.800 atm = 139.463564344 Liters

Answer: volume of gas inside this balloon = 139 L ( 3 significant figures)

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")